# Active Learning of Deep Neural Networks via Gradient-Free Cutting Planes

### v1

@@ F.3 Active Learning Baselines
Finally, we would like to again emphasize that towards a fair and robust comparison, we select the best performing number of epochs (for a discussion on this, refer to Appendix F.4) and learning rate for AL baselines using deep neural networks and use enhancements such as bagging regressors for the scikit-activeml baselines trained on its default regressors.
F.4 Synthetic Spiral Experiment
@@
We now give the deferred supplementary experiment results in Section 8.1. To start, per discussion in Appendix F.2, we would like to discuss the inclusion of regularization through the final convex solve of the equivalent convex program of the two-layer ReLU networks.
Figure 14 shows the final decision boundary with (left) and without (right) the convex solve with the same data and setup as in the spiral experiment in Section 8.1. We abbreviate “after final solve” as AFS and “before final solve” as BFS. It is evident that the inclusion of regularization improves the performance of our cutting-plane active learning method, increasing the accuracy rate on train/test set from 0.84/0.60 to 1.00/1.00. This is therefore the method we use for our cutting-plane AL algorithm. We will see that regularization with the final convex solve does not always speed up the convergence of our cutting-plane AL method in the number of queries. It tends to work less optimally when used on relatively simple dataset, as we will see in Appendix F.5.
@@
F.5 Quadratic Regression Experiment
@@
As in the spiral experiment, we set the hyper-parameters for all deep active learning baselines according to Table 2. In the quadratic regression task, we similarly observe improved performance in deep AL baselines with a higher number of training epochs and a lower learning rate of $1\times 10^{-3}$, ensuring that the baselines provide a robust comparison. For baselines from scikit-activeml, we use the bagging regressor along with the default regressor to enhance their performance (see Appendix F.3). For our cutting-plane AL method (Algorithm 6) and the linear cutting-plane AL method (Algorithm 7), we choose the threshold value $\epsilon=1\times 10^{-3}$.
@@
F.6 IMDB Data Examples

### v3

@@ G.3 Active Learning Baselines
Finally, we would like to again emphasize that towards a fair and robust comparison, we select the best performing number of epochs (for a discussion on this, refer to Appendix G.4) and learning rate for AL baselines using deep neural networks and use enhancements such as bagging regressors for the scikit-activeml baselines trained on its default regressors.
G.4 Synthetic Spiral Experiment
@@
We now give the deferred supplementary experiment results in Section 7.1. To start, per discussion in Appendix G.2, we would like to discuss the inclusion of regularization through the final convex solve of the equivalent convex program of the two-layer ReLU networks.
Figure 14 shows the final decision boundary with (left) and without (right) the convex solve with the same data and setup as in the spiral experiment in Section 7.1. We abbreviate “after final solve” as AFS and “before final solve” as BFS. It is evident that the inclusion of regularization improves the performance of our cutting-plane active learning method, increasing the accuracy rate on train/test set from 0.84/0.60 to 1.00/1.00. This is therefore the method we use for our cutting-plane AL algorithm. We will see that regularization with the final convex solve does not always speed up the convergence of our cutting-plane AL method in the number of queries. It tends to work less optimally when used on relatively simple dataset, as we will see in Appendix G.5.
@@
G.5 Quadratic Regression Experiment
@@
As in the spiral experiment, we set the hyper-parameters for all deep active learning baselines according to Table 2. In the quadratic regression task, we similarly observe improved performance in deep AL baselines with a higher number of training epochs and a lower learning rate of $1\times 10^{-3}$, ensuring that the baselines provide a robust comparison. For baselines from scikit-activeml, we use the bagging regressor along with the default regressor to enhance their performance (see Appendix G.3). For our cutting-plane AL method (Algorithm 6) and the linear cutting-plane AL method (Algorithm 7), we choose the threshold value $\epsilon=1\times 10^{-3}$.
@@
G.6 IMDB Data Examples